1.     The food we eat is degraded, or broken down, in our bodies to provide energy for growth and function. A general overall equation for this very complex process represents the degradation of glucose (C6H12O6) to carbon dioxide (CO2) and water (H2O):

C6H12O6 + 6O2 → 6CO2 + 6H2O

If 968 g of C6H12O6 is consumed by a person over a certain period, what is the mass of CO2 produced?                                                                                         

 

2.     Urea [(NH2)2CO] is prepared by reacting ammonia with carbon dioxide:

2NH3(g) + CO2(g) → (NH2)2CO(aq)+ H2O(l)

In one process, 849.2 g of NH3 are treated with 1223 g of CO2.                       

(a) Which of the two reactants is the limiting reagent?

(b) Calculate the mass of (NH2)2CO formed.

 (c) How much excess reagent (in grams) is left at the end of the reaction?

 

 


1
Expert's answer
2021-12-23T19:28:01-0500

"m(C_6H_{12} O_6) = 968g"


"M(C_6H_{12} O_6) = (12\u00d76)+(1\u00d712)+(16\u00d76) =180g\/mol"


"n(C_6H_{12} O_6)=\frac{m}{M} =\frac{968g}{180g/mol} =5.378mol"


"n(CO_2) =6\u00d75.378 =32.268 mol"


"M(CO_2) =12+16\u00d72 =44g\/mol"


"m(CO_2)=n\u00d7M=32.268mol\u00d744g\/mol =1419.792g"


2.



"n(NH3)=m\/M=849.2g\/17=49.953 mol""n(CO2)=1223g\/44=27.795 mol"

The limiting reagent is NH3. According to the reaction from 49.953 mol NH3 we can get 49.953/2=24.977 mol of (NH2)2CO.



"m((NH2)2CO)=n\u00d7M=24.977\u00d760=1498.62g"


27.795 - 24.977 = 2.818 mol CO2 is left

at the end of the reaction.



"m(CO2)=2.818\u00d744=123.992 g"